AO Commissioning at RFBG: 27 November 2010 UTC

Observer(s): SEsposito, APuglisi, LBusoni, EPinna, DMiller, GBrusa (LBTO)

Visitors(s): RGreen, JBrynnel (LBTO)

Telescope Operator: DHuerta (LBTO)

SW Support: TLiebold (Tucson)

Telescope Support: JLittle (LBTO)

Instrument Support: JMorris (LBTO)

Summary

Preparations

28:30 John filled LBC and MODS

Details

00:00 David opens the telescope, sets IE=-5 CA=47

00:40 Collimate on BS9188

01:00 David checks a light glint in the Chamber. Must have been Safford

01:05 IRAF preset AO816 ACAUIRE gs=21 pam- David sets IE=-11 CA= 50

01:06 IRAF preset AO816 GUIDE gs=21 pa=0

01:07 IRAF preset AO816 ACTIVE gs=21 pa=0

01:08 preset, AO816, ACE, GS=21, RS=0, PA=0 Success

01:09 irc_runao Failed TT modulation

01:13 preset, AO816, ACE, GS=21, RS=0, PA=0 Success

01:15 irc_runao Failed TT modulation

01:20 AO guys work at their Engineering GUI

01:35 Find problem and close loop manually

02:15 preset, AO54, ACE, GS=5, RS=0, PA=180 Success

02:17 Collimate three times

02:20 irc_runao Success

03:03 preset, AO37, ACE, GS=1, RS=0, PA=180 Success

03:07 Collimate three times

03:10 irc_runao Success

03:03 preset, AO892, ACE, GS=7, RS=0, PA=180 Failed Guiding not stable in six iterations. Image has about 3" of astigmatism and the centroid was bouncing from one end of the sausage to the other. Just try to send preset again

03:03 preset, AO37, ACE, GS=1, RS=0, PA=180 Success

03:07 Collimate five times

03:37 irc_runao

04:45 preset, AO1070, ACTIVE, GS=0, RS=0, PA=0 Success

04: Collimate five times

04:45 preset, AO1070, ACE, GS=1, RS=0, PA=0 Success, but no Reference star found on AO TV

04:45 preset, AO107, ACTIVE, GS=0, RS=0, PA=0 Success

03:07 Collimate five times

04:45 preset, AO1070, ACE, GS=1, RS=0, PA=0 Success, but no Reference star found on AO TV

05:45 Point the telescope to horizon, close the shutter and the AO guys put a neutral density in front of the W-Unit

06:25 preset, AO245, ACQUIRE, GS=0 David sets IE=-2 CA=41. David also unwraps the telescope cable chain.

06:30 preset, AO245, ACE, GS=4, RS=0, PA=0 Success

06:47 irc_runao Failed because of timeout in AO Supervisor, but loop is closed. Turn off guiding and Active Optics by hand.

07:40 preset, AO836, ACQUIRE, GS=0 David sets IE=-15 CA=35.

07:43 preset, AO836, ACE, GS=5, RS=0, PA=0 Success

07:44 Collimate three times

07:46 irc_runao Failed Unknown exception

07:47 See active optics is

07:49 David cleared C00

07:57 Run collimate_multiple, 10, first image in /OldRepository/AGW_data/20101123 003. Diverged badly. Stopped.

08:04 preset AO836, ACE, GS=5, RS=0 Improve focus but diverge in other modes. Final wfsc 593.

08:10 preset acquire BS9107 center star

08:15 preset active BS9107 Improve focus but diverge in other modes. Clear C00 and iterate 6 or so times so gs is 1" across

08:30 preset, AO836, GS=5, RS=0, ACE, Success

08:40 runao failed, but close by hand. GS=5 is a binary. This could explain part of it, but BS9107 is definitely not a binary. Don't understand this problem...

09:10 preset, AO911, ACTIVE, GS=0 Success. Collimate for 6-7 iterations. Collimation is not bad too bright

09:15 preset, AO248, ACTIVE, GS=0 Success. Collimate 5 iterations then the primary panice

09:20 preset then runao: AO loop died. While the AO guys were working on this, we power cycle AGW#2

09:30 preset AO248 ACTIVE GS=0 GCS could not get an image from the wfsc.

wfsc 790 now has pupil in the correct place, compared to before we restarted the AGW

11:20 Tested telescope offsets. The functionality is there, but the star (ie the pyramid stages) did not move in the correct direction. We can test this during the day.

11:35 sent a preset to position the star on the IRTC. We rotated the instrument rotator 180 and saw the star, the was positioned on the bottom center edge of the IRTC image, rotated around the center of the IRTC frame. Thus, in the case of this star, either the target or the off-axis guide star coordinates are not correct.

11:50 Sent a preset AO248 ACE, GS=2, RS=0. Alfio change this to a TTM preset on the AO Engineering GUI. A runao then closed the tip/tilt field stabilization loop. They collected some data.

12:34 preset AO301 TTM GS=2 RS=0 Failed because the reference star was too bright (6.35, where max is 8.5). I modified AO301.catalog

12:35 preset AO301 TTM GS=2 RS=0 Success

12:36 runao Success

12:50 preset OC_NGC2281, GS=10, RS=1, ACE. Saw reference star on AO TV, saw target star on IRTC (but not Reference star) and guide star on guider. However, the Reference star was being vignetted by the Neutral Desity. It is getting bright outside so call it a night.

-- DougMiller - 27 Nov 2010
Topic revision: r4 - 27 Nov 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback